OpenStudy (anonymous):

Find the number of terms in these geometric series 0.03+0.06+0.12+....+1.92

OpenStudy (anonymous):

In a geometric series the n-th term is \(T_n = ar^{n-1} \)

OpenStudy (anonymous):

\[.03(1+2+2^2+2^3+...+2^6)=.03\times( 2^7-1)\]

OpenStudy (anonymous):

\[1.92 = .03 \times 2^{n-1} \implies 2^6 = 2^{n-1} \implies n = 7 \]